# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ Appendix B PROOF OF MAIN TEXT RESULTS
and
$\mu_{\widehat{Y}|\hat{z}_{1}}=\mu_{\widehat{Y}}+\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}\bar{\bm{\tau}}_{1}\left(\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}\bar{\bm{\tau}}_{1}\right)^{-1}\left(\hat{z}_{1}-\mu_{Z_{1}}\right)$
Ṫhus using the above equation
$\mu_{\widehat{Y}|\hat{z}_{1}}=\mu_{\widehat{Y}}+\alpha\left(\hat{z}_{1}-\mu_{Z_{1}}\right)=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\bm{\mu}_{\mathbf{U}_{0}}+\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}\bm{\mu}_{\pi(1)}+\alpha\hat{z}_{1}-\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\bm{\mu}_{\pi(1)}\,,$
@@
$\mbox{var}({\widehat{Y}|\hat{z}_{1}})=\sigma_{\widehat{Y}}^{2}-\bm{\tau}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{01}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}\bar{\bm{\tau}}_{1}\left(\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}\bar{\bm{\tau}}_{1}\right)^{-1}\\ \bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}A_{01}^{\top}(I_{N_{0}}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}$
@@
$\mbox{var}({\widehat{Y}|\hat{z}_{1}})=\sigma_{\widehat{Y}}^{2}-\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}A_{01}^{\top}(I_{N_{0}}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}\\ =\sigma_{\widehat{Y}}^{2}-\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{11})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{11})^{-\top}\bar{\bm{\tau}}_{1}\alpha\\ =\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\Sigma_{\pi(0)}(I_{N_{0}}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}\,.$
@@
where we can identify all parameters with the above equations.
B.3 Proof of Proposition 3.4
See 3.4
Proof.
such that $\mathbf{X}$, $A$ and $P$ take the block forms
with $\pi(0)$ of size $N_{0}$, and $\pi(1)$ of size $N_{1}=N-N_{0}$ and $S$ of size $s$. Then we know from the Proposition 3.3 that there is already a valid solution using $\pi$ as alignment. The only difference is that variables in $S$ are unintervened, which does not affect the ability of the solution to achieve $\mathcal{L}_{cons}=0$. That solution would be compatible with interventions on $S$, but since $S$ is unintervened, we do not have uniqueness guarantees for this choice of $\pi$.
Alternatively, if we choose $\pi^{\prime}(0)=\pi(0)\cup S$ and $\pi^{\prime}(1)=\pi(1)\setminus S=\Omega$, then, we can again apply Propostion 3.4, and see that it provides a different solution with this alignment, which is compatible with the given problem (constructive transformation with constraint on the mapping $\tau_{0}$). Importantly, the key indeterminacy is for the map $\tau_{1}$, which will either put all its weight on elements in $S$ (direct parents of $\pi(0)$), or alternatively, put all its weights on elements in $\Omega$. There is an additional, but trivial, indeterminacy for the map $\omega_{1}$: indeed, since $X_{S}$ is unintervened (as part of $\pi(0)$), the weights in ${\bm{\omega}}_{1}$ associated to these coefficients may take arbitrary values (since their associated component in $\mathbf{i}$ remains zero). We do not consider these trivial indeterminacies (which do not affect the mapping $\omega_{1}$ on its domain, i.e. the support of the prior $P(\mathbf{i})$) by forcing the weights of ${\bm{\omega}}_{1}$ associated to unintervened variables to zero. ∎
Appendix C THE CASE $n>1$: MULTIPLE HIGH-LEVEL CAUSES
C.1 Updated setting
We generalize the main text setting of the linear reduction framework to learn $n>1$ high-level causes as follows.
Tau map
The $\tau$-map now incorporates $n+1$ components, associated to each cause and to $Y$, respectively.
$\mathbf{X}\mapsto\begin{bmatrix}Y\\ {\mathbf{Z}}\end{bmatrix}=\begin{bmatrix}\tau_{0}(\mathbf{X}_{\pi(0)})\\ T\mathbf{X}\end{bmatrix}=\begin{bmatrix}\bm{\tau}_{0}^{\top}\mathbf{X}_{\pi(0)}&\bm{\tau}_{1}^{\top}\mathbf{X}_{\pi(1)}&\dots&\bm{\tau}_{n}^{\top}\mathbf{X}_{\pi(n)}\end{bmatrix}^{\top}$
Omega map
We define the collection of $n$ vectors $\{{\bm{\omega}}_{k}\}$ associated to interventions on each high-level cause. Again, no intervention is performed on the target $Y$.
$\omega(\mathbf{i})=W\mathbf{i}=[{\bm{\omega}}_{1}^{\top}\bm{i},\,...,\,{\bm{\omega}}_{n}^{\top}\bm{i}]^{\top}=[\bar{{\bm{\omega}}}_{1}^{\top}\bm{i}_{\pi(1)},\,...,\,\bar{{\bm{\omega}}}_{n}^{\top}\bm{i}_{\pi(n)}]^{\top}\,,$
High-level mechanisms
As illustrated by Fig.0(b), we now have $n$ causes are root nodes, which implies that the high-level model represents their mechanism as sampling from independent scalar exogenous variables. In the Gaussian case, this entails modeling their joint distribution by a multivariate Gaussian with diagonal covariance. The additional mechanism, associated with $Y$, is now a function $f_{\bm{\gamma}}$ that simply takes multiples causes as input, in addition to its independent exogenous variable $R_{0}$. In the linear Gaussian SCM case, this is simply implemented by applying multiplicative coefficients to each input and summing the outcome.
C.2 Additional indeterminacies and group sparsity loss
In the $n=1$ high-level cause setting, choosing $\pi(1)$ is unproblematic, as one can just select variables relevant for interpretability among all components of $\mathbf{X}$ that are not used to compute $Y$. In contrast, for $n>1$ causes, the choice of the subsets $\pi(1),\dots,\pi(n)$ is non-trivial. One the one-hand, being strict about imposing constructive transformations would entail that those subsets do not have any pairwise overlap. On the other hand, causes “compete” to account for the interventional changes in $Y$, which may lead them to put large weights on the same variables in their vectors $\bm{\tau}_{k}$ and ${\bm{\omega}}_{k}$. Additionally, there is nothing in the consistency loss preventing one single cause, say $Z_{1}$, taking over all the causal explanation of $Y$, while the other high-level causes $Z_{2},\dots,Z_{n}$ may just be chosen as irrelevant variables independent of (and having no influence on) $Y$.
In order to prevent a single cause taking over and softly enforce that each of them focus on different subsets of low-level variables, we take inspiration from losses enforcing structured sparsity (Jenatton et al., 2011). We assume that domain-knowledge can inform us about subsets of low-level variables, preferably belong to the same subset $\pi(k)$. The idea behind is to group variables of the same nature, pertaining to the same mechanism, or associated to neighboring location in the coordinate system. In this way, the cause $Z_{k}$ associated to subset $\pi(k)$ may be interpreted as conveying information about an interpretable aspect of the model. Let $G_{1},\ldots,G_{M}\subseteq[1..N]$ define such groups of low-level variables. Then, we define the group sparsity loss as
$\sum_{j=1}^{n}\frac{\left(\sum_{m=1}^{M}\sqrt{\sum_{k\in G_{m}}(\bm{\tau}_{j})_{k}^{2}}\right)^{2}}{\sum_{m=1}^{M}\sum_{k\in G_{m}}(\bm{\tau}_{j})_{k}^{2}}-n.$ (18)
While the numerator corresponds to the family of convex group sparsity inducing losses (Jenatton et al., 2011), we adapt it to our setting by normalizing it by an $\ell_{2}$ squared norm at the numerator in order to make this loss invariant by multiplicative rescaling of the $\bm{\tau}_{k}$ vectors. This avoids the algorithm to downscale the reduction map This regularization encourages the different high-level variables to focus on different groups of nodes and is a way to encode prior knowledge about the system in the TCR. The use of this loss is exemplified in App. F.2 and Fig. 9.
@@ D.3 The case of linear chain SCMs
Transposing the chain example to the case of Proposition 3.4, we can take the case were the direct parent $X_{N-1}$ of the target is left unintervened. In such a case, $\bar{\bm{\tau}}_{1}$ may put its weight on both $X_{N-1}$ and its direct parent $X_{N-2}$, Proposition 3.3 provides two example solutions for different choices of $\pi(1)$, including or excluding $X_{N-1}$. In the most extreme case of dissimilarity between $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$, solution including $X_{N-1}$ in $\pi(1)$ puts all $\bm{\tau}_{1}$’s weight on $X_{N-1}$, while ${\bm{\omega}}_{1}$ has no weight on it (because it is unintervened). As a consequence, ${\bm{\omega}}_{1}$ and $\bm{\tau}_{1}$ are orthogonal and the associated homogeneity loss vanishes. In contrast, the unique solution excluding $X_{N-1}$ from $\pi(1)$ have a larger cosine similarity and will thus be preferred by the homogenity-regularized loss.

### v2

@@ Appendix C PROOF OF MAIN TEXT RESULTS
and we can derive its conditional mean and covariance
$\mu_{\widehat{Y}|\hat{z}_{1}}=\mu_{\widehat{Y}}+\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1}\left(\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1}\right)^{-1}\left(\hat{z}_{1}-\mu_{Z_{1}}\right)\,.$
Thus using the above equation
$\mu_{\widehat{Y}|\hat{z}_{1}}=\mu_{\widehat{Y}}+\alpha\left(\hat{z}_{1}-\mu_{Z_{1}}\right)=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\bm{\mu}_{\pi(0)}+\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\bm{\mu}_{\Omega}+\alpha\hat{z}_{1}-\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\bm{\mu}_{\Omega}\,,$
@@
$\mbox{var}({\widehat{Y}|\hat{z}_{1}})=\sigma_{\widehat{Y}}^{2}-\bm{\tau}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{0\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1}\left(\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1}\right)^{-1}\\ \bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\Omega}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}A_{0\Omega}^{\top}(I_{N_{0}}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}$
@@
$\mbox{var}({\widehat{Y}|\hat{z}_{1}})=\sigma_{\widehat{Y}}^{2}-\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}A_{0\Omega}^{\top}(I_{N_{0}}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}\\ =\sigma_{\widehat{Y}}^{2}-\alpha\bar{\bm{\tau}}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\Sigma_{\pi(1)}(I_{N_{1}}-A_{\Omega\Omega})^{-\top}\bar{\bm{\tau}}_{1}\alpha\\ =\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\Sigma_{\pi(0)}(I_{N_{0}}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}\,.$
@@
where we can identify all parameters, including the mean and variance of the Gaussian exogenous variable $R_{0}$, with the above Eqs (22,24,25).
Part 2: $n$-causes case ($n>1$)
We now assume ${(\tau_{k}^{\prime},\omega_{k}^{\prime},\pi^{\prime})}_{k=1..n}$ such that the loss vanishes, but no such solution for $n+1$ causes.
Properties of exact $n$-cause solutions Then such a solution can be linked to the 1-cause solution, which is guaranteed to exist according to our set of assumptions. Indeed, the existence of the n-cause solution implies that the pushfoward interventional distribution of the low-level causal model by $\tau$, $\widehat{P}$ satisfies
$\widehat{P}^{(\mathbf{i})}(Y|\mathbf{Z}=\mathbf{z})\sim\mathcal{N}\left(\sum_{k}\alpha_{k}z_{k}+\beta,\sigma_{Y|Z}^{2}\right)$
and
$\widehat{P}^{(\omega(\mathbf{i}))}(\mathbf{Z})\sim\prod_{k}\widehat{P}^{(0)}(Z_{k}-\omega_{k}(\mathbf{i}_{k}))$
If we define the aggregate cause $\tilde{Z}=\sum_{k}\alpha_{k}z_{k}=\sum_{k}\alpha_{k}\tau_{k}(x)$, then we can rewrite the above model as
$\widehat{P}^{(\mathbf{i})}(Y|\tilde{Z}=\tilde{z})\sim\mathcal{N}\left(\tilde{z}+\beta,\sigma_{Y|Z}^{2}\right)$
and
$\widehat{P}^{(\omega(\mathbf{i}))}(\tilde{Z})\sim\prod_{k}\widehat{P}^{(0)}(\tilde{Z}-\sum_{k}\omega_{k}(\mathbf{i}_{k}))$
which implies that concatenating the $\tau_{k}$ with multiplicative coefficient $\alpha_{k}$ leads to a valid 1-cause TCR, and must thus match the expressions we have found for it, up to a multiplicative constant.
Moreover, the interventional consistency of the $n$-causes, which do not influence each other according to the assumed high-level causal graph, entails that any low-level intervention $\mathbf{i}$ affects only high-level variable $Z_{k}$ through is components in $\pi(k)$.
We define $A_{1..n,1..n}$ by reordering the indices of $\Omega$ according to the assumed alignment $\pi$. Consistency then implies (using the above 1 cause solution proof)
$\begin{bmatrix}\bar{\bm{\tau}}_{1}^{\top},&\dots,&0\\ \vdots,&\ddots,&\vdots\\ 0,&\dots,&\bar{\bm{\tau}}_{n}^{\top}\end{bmatrix}(I_{\sum_{k}N_{k}}-A_{1..n,1..n})^{-1}\mathbf{i}_{\Omega}=\begin{bmatrix}{\bm{\omega}}_{1}^{\top}\mathbf{i}_{\pi_{1}}\\ \vdots\\ {\bm{\omega}}_{n}^{\top}\mathbf{i}_{\pi_{n}}\end{bmatrix}\,.$ (26)
This implies that
$\bar{\bm{\tau}}_{k}^{\top}(I_{\sum_{k}N_{k}}-A_{1..n,1..n})^{-1}_{kj}=0\,\mbox{ for all }j\in\pi(l),l\neq k\,,$ (27)
where (.$(.)_{kj}$ indicates the matrix block corresponding to indices in $\pi(k)\times\pi(j)$. Because the non-vanishing coefficients of $\bar{\tau_{k}}$ reflect the influences along the causal pathways from nodes of $\pi(k)$ to $Y$, the above entails that $(I_{\sum_{k}N_{k}}-A_{1..n,1..n})^{-1}_{kj}$ must vanish on the support of $\bar{\tau_{k}}$. Indeed, otherwise Eq. (27) would indicate that causal pathways from nodes in $\pi(k)$ to $\pi(0)$ cancel each other, which is forbidden by our assumptions.
We thus deduce that any off-diagonal block element of $(I_{\sum_{k}N_{k}}-A_{1..n,1..n})^{-1}$ whose row component belongs to the support of any $\tau_{k}$ and whose column component belongs to the support of any $\omega_{k}$ must vanish. Indeed, otherwise the causes would influence each other. In essence, this means that any node influencing the target must not influence any node though some causal pathway in another group with the same property.
Identifiability of the $n$-cause solution We assume $n=n_{max}$ and consider a second $n$-cause solution, which we denote: ${(\tau_{k}^{\prime},\omega_{k}^{\prime},\pi^{\prime})}_{k=1..n}$.
If $\pi^{\prime}=\pi$ up to a permutation of the order of the causes and removal of low-level variables that do not belong to the support of neither any $\omega$. Then identifiability of the corresponding 1-cause solution implies that each $\tau_{k}^{\prime}$ is identified with each $\tau_{k}$ up to a multiplicative constant, because they both match the components of the 1-cause $\tau$ on their (identical) support $\pi(k)$ The same goes for $\omega_{k}^{\prime}$ and $\omega_{k}$. This corresponds to the conclusion of the Proposition.
Otherwise, $\pi^{\prime}\neq\pi$ even up to a permutation of the order of the causes and removal of low-level variables that do not belong to the support of neither any $\tau$ nor $\omega$. There should be an overlap between supports of omegas and taus of one cause of one solution with two different causes of the other solution. Without loss of generality, because of the block-diagonal structure of $(I_{\sum_{k}N_{k}}-A_{1..n,1..n})^{-1}$ entailed by Eqs. (26-27) for both solutions, this overlap implies that $\pi(k)$ for at least one $k$ can be further partitioned and reordered into two subgroups, such that the corresponding diagonal block of $(I_{\sum_{k}N_{k}}-A_{1..n,1..n})^{-1}$ can be turned in a block diagonal submatrix. This can be used to build a new alignment $\pi^{\prime\prime}$ for $n+1$ causes, and its associated tau and omega maps such that Eq. (26) will be again satisfied, leading to interventional consistency of the (n+1)-causes. Moreover, because the exogenous variables in $\Omega$ are assumed independent, the block diagonal structure of the newly defined matrix $(I_{\sum_{k}N_{k}}-A_{1..n+1,1..n+1})^{-1}$ entails that the covariance of the $n+1$ high-level cause variable will be diagonal, ensuring mutual independence between high-level causes. This procedure exhibits the existence of an $(n+1)$-cause TCR, contradicting the original assumption that $n=n_{max}$. This case is thus excluded. ∎
@@ D.3 The case of linear chain SCMs
Transposing the chain example to the case of Proposition D.2, we can take the case were the direct parent $X_{N-1}$ of the target is left unintervened. In such a case, $\bar{\bm{\tau}}_{1}$ may put its weight on both $X_{N-1}$ and its direct parent $X_{N-2}$, Proposition 3.4 provides two example solutions for different choices of $\pi(1)$, including or excluding $X_{N-1}$. In the most extreme case of dissimilarity between $\bm{\tau}_{1}$ and ${\bm{\omega}}_{1}$, solution including $X_{N-1}$ in $\pi(1)$ puts all $\bm{\tau}_{1}$’s weight on $X_{N-1}$, while ${\bm{\omega}}_{1}$ has no weight on it (because it is unintervened). As a consequence, ${\bm{\omega}}_{1}$ and $\bm{\tau}_{1}$ are orthogonal and the associated homogeneity loss vanishes. In contrast, the unique solution excluding $X_{N-1}$ from $\pi(1)$ have a larger cosine similarity and will thus be preferred by the homogenity-regularized loss.
D.4 Loss of identifiability through unintervened variables
Figure 5: 1-cause TCR solutions on a chain graph. Arrows indicate non-zero coefficients of each map. (a) Unique solution $\bm{\tau}_{1}$ when interventions are performed on all nodes except the target. (b) Two solutions $\bm{\tau}_{1}$ and $\bm{\tau}^{\prime}_{1}$ when only the first two nodes are intervened on.
Proposition D.2.
Consider the setting of Prop. 3.4 with the exception that $\Omega\subsetneqq\pi(1)$ such that there is now a non-empty subset $S=\pi(1)\setminus\Omega$, such that $\mathbf{X}_{\Omega}\to\mathbf{X}_{S}\to\mathbf{X}_{\pi(0)}$. Then there exist at least two different linear 1D TCR such that $\mathcal{L}_{\mathrm{cons}}=0$.
This result can also be illustrated with a chain graph, as shown in Fig. 5(b). If the parent node $X_{3}$ of $Y=X_{4}$ is unintervened, then one may choose either $Z_{1}=X_{2}$ or $Z_{1}=X_{3}$ (matching the solution of Fig. 5(a)) to minimize $\mathcal{L}_{\mathrm{cons}}$. This is because both variables are equivalently mediating all performed interventions to $X_{4}$. Note that each choice has it own benefit: $Z_{1}=X_{3}$, as a direct parent of $Y$, is a better statistical predictor of the value of $Y$. However, if we focus on causal interpretability of the high-level representation, $Z_{1}=X_{2}$ is preferable because it is one of the variables intervened on at the low-level as enforced by the prior $P(\mathbf{i})$, and such that it will be associated to a non-zero weight in ${\bm{\omega}}_{1}$ for any solution satisfying $\mathcal{L}_{\mathrm{cons}}=0$.
Proof.
such that $\mathbf{X}$, $A$ and $P$ take the block forms
with $\pi(0)$ of size $N_{0}$, and $\pi(1)$ of size $N_{1}=N-N_{0}$ and $S$ of size $s$. Then we know from the Proposition 3.4 that there is already a valid solution using $\pi$ as alignment. The only difference is that variables in $S$ are unintervened, which does not affect the ability of the solution to achieve $\mathcal{L}_{\mathrm{cons}}=0$. That solution would be compatible with interventions on $S$, but since $S$ is unintervened, we do not have uniqueness guarantees for this choice of $\pi$.
Alternatively, if we choose $\pi^{\prime}(0)=\pi(0)\cup S$ and $\pi^{\prime}(1)=\pi(1)\setminus S=\Omega$, then, we can again apply Propostion D.2, and see that it provides a different solution with this alignment, which is compatible with the given problem (constructive transformation with constraint on the mapping $\tau_{0}$). Importantly, the key indeterminacy is for the map $\tau_{1}$, which will either put all its weight on elements in $S$ (direct parents of $\pi(0)$), or alternatively, put all its weights on elements in $\Omega$. There is an additional, but trivial, indeterminacy for the map $\omega_{1}$: indeed, since $X_{S}$ is unintervened (as part of $\pi(0)$), the weights in ${\bm{\omega}}_{1}$ associated to these coefficients may take arbitrary values (since their associated component in $\mathbf{i}$ remains zero). We do not consider these trivial indeterminacies (which do not affect the mapping $\omega_{1}$ on its domain, i.e. the support of the prior $P(\mathbf{i})$) by forcing the weights of ${\bm{\omega}}_{1}$ associated to unintervened variables to zero. ∎
D.5 Connection to causal abstractions
Proposition D.3.
Assume the low-level SCM follows
$\mathbf{X}\coloneqq A\mathbf{X}+\mathbf{U}+\mathbf{i}\,,\quad U_{k}\sim\mathcal{N}(\mu_{k},\sigma_{k}^{2})\,,\sigma_{k}^{2}>0\,,\,\,\mathbf{i}\sim P(\mathbf{i})$
such that $\mathbf{X}$ and $A$ take the block forms
Given an arbitrary choice of linear scalar target of the form $Y=\tau_{0}^{\top}\mathbf{X}=\bar{\tau}_{0}^{\top}\mathbf{X}_{\pi(0)}$, under the conditions of Proposition 3.4, the unique linear 1-cause TCR (up to a multiplicative constant) is associated to a 1-cause constructive abstraction given by
$\bar{\bm{\tau}}_{1}=$ $\,A_{01}^{\top}(I_{\#\pi(0)}-A_{00})^{-\top}\bar{\bm{\tau}}_{0}$ (34)
$\quad\bar{{\bm{\omega}}}_{1}=$ $\,(I_{\#\pi(1)}-A_{11})^{-\top}\bar{\bm{\tau}}_{1}\,,$ (35)
$\bar{\bm{\tau}}_{U,0}=$ $\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\,,$ (36)
and $\text{and }\bar{\bm{\tau}}_{U,1}=$ $\bar{\tau}_{1}^{\top}(I_{N_{1}}-)^{-1}\,.$ (37)
Proof.
To have a valid constructive causal abstraction, we need to verify the existence of an additional constructive map $\tau_{U}$ for exogenous variables such that for all realizations $\mathbf{u}$ of $\mathbf{U}$.
$\tau(\mathcal{L}^{(\mathbf{i})}(\mathbf{u}))=\mathcal{H}^{(\omega(\mathbf{i}))}(\tau_{U}(\mathbf{u}))$
where $\mathcal{L}^{(\mathbf{i})}(.)$ and $\mathcal{H}^{(\mathbf{j})}(.)$ denote the mappings from endogenous to exogenous variable for the low- and high-level intervened models, respectively.
Using
$\mathcal{L}^{(\mathbf{i})}(\mathbf{U})=\mathbf{X}^{(\mathbf{i})}=(I_{N}-A)^{-1}(\mathbf{U}+\mathbf{i})$
we get
$\mathbf{X}_{\pi(1)}^{(\mathbf{i})}=(I_{N_{1}}-)^{-1}(\mathbf{U}_{\pi(1)}+\mathbf{i}_{\pi(1)})$
and
$\mathbf{X}_{\pi(0)}^{(\mathbf{i})}=(I_{N_{0}}-A_{00})^{-1}\mathbf{U}_{\pi(0)}+(I_{N_{0}}-A_{00})^{-1}A_{01}\mathbf{X}_{\pi(1)}^{(\mathbf{i})}\,.$
Applying the $\tau_{0}$ map we get the solution of Prop. 3.4
$Y^{(\mathbf{i})}=\bar{\bm{\tau}}_{0}^{\top}\mathbf{X}_{\pi(0)}^{(\mathbf{i})}=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\mathbf{U}_{\pi(0)}+\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}A_{01}\mathbf{X}_{\pi(1)}^{(\mathbf{i})}=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}\mathbf{U}_{\pi(0)}+\bar{\bm{\tau}}_{1}^{\top}\mathbf{X}_{\pi(1)}^{(\mathbf{i})}$
moreover applying the ${\tau}_{1}$ map to the variables in $\pi(1)$ we get
$\bar{\tau}_{1}^{\top}\mathbf{X}_{\pi(1)}^{(\mathbf{i})}=\bar{\tau}_{1}^{\top}(I_{N_{1}}-)^{-1}(\mathbf{U}_{\pi(1)}+\mathbf{i}_{\pi(1)})=\bar{\tau}_{1}^{\top}(I_{N_{1}}-A_{\Omega\Omega})^{-1}\mathbf{U}_{\pi(1)}+\bar{{\bm{\omega}}}_{1}^{\top}\mathbf{i}_{\pi(1)}$
So by defining the vectors $\bar{\bm{\tau}}_{U,1}^{\top}=\bar{\tau}_{1}^{\top}(I_{N_{1}}-)^{-1}$ and $\bar{\bm{\tau}}_{U,0}^{\top}=\bar{\bm{\tau}}_{0}^{\top}(I_{N_{0}}-A_{00})^{-1}$ we get a valid constructive abstraction linking the low-level map $\mathbf{u}\mapsto\mathcal{L}^{(\mathbf{i})}(\mathbf{u})$ to the following high-level map (first component is $Y$, second is the cause $Z_{1}$)
$\mathbf{r}\mapsto\mathcal{H}^{(\mathbf{j})}(\mathbf{r})=\begin{bmatrix}r_{0}+r_{1}+j_{1}\\ r_{1}+j_{1}\end{bmatrix}$
such that for all $\mathbf{u}$
$\mathcal{H}^{(\bar{{\bm{\omega}}}_{1}^{\top}\mathbf{i}_{\pi(1)})}\left(\begin{bmatrix}\bar{\bm{\tau}}_{U,0}^{\top}\mathbf{u}_{\pi(0)}\\ \bar{\bm{\tau}}_{U,1}^{\top}\mathbf{u}_{\pi(1)}\end{bmatrix}\right)=\tau(\mathcal{L}^{(\mathbf{i})}(\mathbf{u}))$